15 years from now, Johnny's age will be twice his age 10 years ago. What is his current age?
Thank you for the assistance!

Answer by macston(5194)

Johnny's age=J
J+15=2(J-10)
J+15=2J-20 add 20 to both sides
J+15+20=2J-20+20
J+35=2J Subtract J from each side
J+35-J=2J-J
35=J ANSWER Johnny is 35 years old
CHECK:
15 years from now Johnny is 50, which is twice his age 10 years ago(35-10=25 years)